OpenStudy (anonymous):

Just making sure, the correct answer is 12, right? Problem- ABCD is a parallelogram. If AB is equal to 2x + 1 and CD is equal to x + 13, find x.

OpenStudy (confluxepic):

Yes the answer is 12.

OpenStudy (confluxepic):

2x+1=x+13 Subtract 1. 2x=x+12 Subtract x. x=12